Compute the monthly payments on a 3-year lease for a $25,704 car if the annual rate of depreciation is 14% and the lease's annual interest rate is 5.5%. Round your answer to the nearest dollar.
Answer #1

Monthly payment = car amount / cummulative discount factor

Cummulative discount factor = (1-(1+i)^-n)/i

I= 5.5% /12 = 0.0045833

n= 3*12= 36

Cummulative discount factor = (1-(1+0.004583)^-36)/0.0045833= 33.1173

Month payment = 25704/ 33.1173= 776.15